OpenStudy (anonymous):

Give exact & approximate solutions to three decimal places (x-9)^2=4

OpenStudy (chaise):

Expand your brackets: (x-9)^2=4 x^2-18x+81=4 Bring the 4 to the other side of the equation, making it equal to 0. x^2-18x+77=0 Factor: (x-11)(x-7) x-11=0 x=11 or x-7 = 0 x =7